### 14 Mar 2024 v2

# On closed definable subsets in Hensel minimal structures

Abstract.
This paper deals with Hensel minimal structures on non-trivially valued fields $K$. The main aim is to establish the following two properties of closed 0-definable subsets $A$ in the affine spaces $K^{n}$. Every such subset $A$ is the zero locus of a continuous 0-definable function $f:K^{n}\to K$, and there exists a 0-definable retraction $r:K^{n}\to A$. While the former property is a non-Archimedean counterpart of the one from o-minimal geometry, the former does not hold in real geometry in general. The proofs make use of a model-theoretic compactness argument and ubiquity of clopen sets in non-Archimedean geometry.
Key words and phrases:
Non-Archimedean geometry, Hensel minimality, closed sets, cell decomposition, definable retractions, definable extension
2000 Mathematics Subject Classification:
Primary 03C65, 03C98, 12J25.
1. Introduction
This paper continues our research on Henselian valued fields from our papers [14, 16, 17, 18], and particularly on Hensel minimal structures from [19, 20]. The axiomatic theory of Hensel minimal structures was introduced by Cluckers–Halupczok–Rideau [4, 5]. Actually, their research followed numerous earlier attempts to find suitable approaches in geometry of Henselian valued fields which, likewise o-minimality in real geometry, would realize the postulates of both tame topology and tame model theory. Those attempts have led to various, axiomatically based concepts such as C-minimality [11, 15], P-minimality [12], V-minimality [13], b-minimality [6], tame structures [2, 3], and eventually Hensel minimality.
So let $K$ be a 1-h-minimal structure, i.e. a model of a 1-h-minimal theory in an expansion $\mathcal{L}$ of the language of valued fields. The main aim is to establish the following two properties of closed 0-definable subsets $A$ in the affine spaces $K^{n}$.
Theorem 1.1.
Every closed 0-definable subset $A$ of $K^{n}$ is the zero locus $\mathcal{Z}(g):=\{x\in K^{n}:\ g(x)=0\}$ of a continuous 0-definable function $g$ on $K^{n}$.
Theorem 1.2.
For every non-empty closed 0-definable subset $A$ of an affine space $K^{n}$, there exists a 0-definable retraction $r:X\to A$.
Remark 1.3.
Note that while the former property is a counterpart of the one from o-minimal geometry (see e.g. [1, Proposition 2.7.5] and [8, Property 4.22]), the former does not hold in real geometry in general. The proofs of these theorems make use of a model-theoretic compactness argument and ubiquity of clopen sets in non-Archimedean geometry.
As an immediate corollary from Theorem 1.2, we obtain the following non-Archimedean version of the Tietze–Urysohn extension theorem.
Corollary 1.4.
Let $A$ be a closed 0-definable subset of an affine space $K^{n}$. Then every continuous 0-definable function $f:A\to K$ can be extended to a continuous 0-definable function $F:X\to K$.
Proof.
Put $F:=f\circ r$ where $r:K^{n}\to A$ is a 0-definable retraction. ∎
We shall adopt the multiplicative convention ||$\cdot$|| ⋅ | for the valuation of the ground field $K$. For $x\in K^{n}$, put
$|x|:=\max\,\{|x_{1}|,\ldots,|x_{n}|\}$
and
$B(a,r):=\{x\in K^{n}:\ |x-a|<r\}$
for $a\in K^{n}$, $r\in|K|$, $r>0$. Note that every two (open) balls $B(a,\epsilon)$ and $B(x,r)$, $a,x\in K^{n}$, are either disjoint, or one is contained in the other according as $\epsilon\leq r$ or $r\leq\epsilon$.
2. Proofs of the main results
We begin by stating two lemmas.
Lemma 2.1.
If $A=\bigcup_{i=1}^{s}A_{i}$ and the conclusions of the main theorems hold for every subset $A_{i}$, then they hold for $A$.
Proof.
The case of Theorem 1.1 is immediate. Indeed, if $A_{i}=\mathcal{Z}(g_{i})$, $i=1,\ldots,s$, then $A=\mathcal{Z}(g_{1}\cdot\ldots\cdot g_{s})$.
For Theorem 1.2, it is enough (by induction on $s$) to consider the case $s=2$. Put
$U_{1}:=\bigcup\,\left\{B(x,r):\ x\in A_{1}\setminus A_{2},\ r<R(x)\right\},$
where
$B(x,r):=\{y\in K^{n}:\ |y-x|<r\}$
and
$R(x):=\{|x-y|:\ y\in A_{2}\}\ \ \text{for}\ \ x\in A_{1}\setminus A_{2}.$
Obviously, $U_{1}$ is an open 0-definable subset of $K^{n}$ containing $A_{1}\setminus A_{2}$ and disjoint from $A_{2}$.
Furthermore, we assert that the set $U_{1}\cup(A_{1}\cap A_{2})$ is a closed subset of $K_{n}$. It suffices to show that if $a$ lies in the closure $\overline{U_{1}}$ of the set $U_{1}$, then $a\in U_{1}$ or $a\in A_{1}\cap A_{2}$. So supposing $a\not\in A_{1}\cap A_{2}$, we must get $a\in U_{1}$.
Were $a\in A_{2}$, we would get $a\not\in A_{1}$. Then
$\forall\,\epsilon>0\ \exists\,x\in A_{1}\setminus A_{2}\ \exists\,r<R(x)\ \ B(a,\epsilon)\cap B(x,r)\neq\emptyset.$
If $\epsilon\leq r$, then $B(a,\epsilon)\subset B(x,r)$, and thus $B(x,r)\cap A_{2}\neq\emptyset$, which is impossible. Therefore $\epsilon\geq r$, and then $B(x,r)\subset B(a,\epsilon)$. Hence $B(a,\epsilon)\cap A_{1}\neq\emptyset$ for all $\epsilon>0$, and thus $a\in\overline{A_{1}}=A_{1}$, which again is a contradiction.
Therefore $a\not\in A_{2}$ and then $B(a,\epsilon_{0})=\emptyset$ for an $\epsilon_{0}>0$. When $a\in A_{1}$, then $a\in U_{1}$, as desired. So suppose $a\not\in A_{1}$. Then
$\forall\,\epsilon\in(0,\epsilon_{0})\ \exists\,x\in A_{1}\setminus A_{2}\ \exists\,r<R(x)\ \ B(a,\epsilon)\cap B(x,r)\neq\emptyset.$
If $\epsilon\leq r$ for some $r<R(x)$, then $B(a,\epsilon)\subset B(x,r)$, $a\in B(x,r)$ and thus $a\in U_{1}$, as desired. The other case where $\epsilon\geq r$ for all $r<R(x)$ is impossible, because then we would get $B(x,r)\subset B(a,\epsilon)$. Hence $B(a,\epsilon)\cap A_{1}\neq\emptyset$ for all $\epsilon>0$, and thus $a\in\overline{A_{1}}=A_{1}$, which again is a contradiction. This proves the assertion.
Consequently, $U_{1}$ is disjoint from $A_{2}$ and is a clopen 0-definable neighbourhood of $A_{1}\setminus A_{2}$ in $K^{n}\setminus(A_{1}\cap A_{2})$.
The situation is symmetric. It is easy to check that the set
$U_{2}:=K^{n}\setminus(U_{1}\cup(A_{1}\cap A_{2})$
is disjoint from $A_{1}$, is a clopen 0-definable neighbourhood of $A_{2}\setminus A_{1}$ in $K^{n}\setminus(A_{1}\cap A_{2})$, and the set $U_{2}\cup(A_{1}\cap A_{2})$ is a closed subset of $K^{n}$.
Therefore, if $r_{1}:X\to A_{1}$ and $r_{2}:X\to A_{2}$ are two 0-definable retractions, so is the map
$r:K^{n}\to(A_{1}\cup A_{2})$
given by the formula
$r(x)=\left\{\begin{array}[]{cl}r_{1}(x)&\mbox{ if }\ x\in U_{1}\cup(A_{1}\cap A_{2}),\\ r_{2}(x)&\mbox{ if }\ x\in U_{2}\cup(A_{1}\cap A_{2}),\end{array}\right.$
which is the desired result. ∎
Lemma 2.2.
There exists a continuous definable function $g_{t}:K_{w}^{t}\to K$ such that
$g_{t}(w)=0\ \ \Longleftrightarrow\ \ w=0.$
Proof.
We define the functions $g_{t}$ inductively. Put
$g_{2}(w_{1},w_{2}):=\left\{\begin{array}[]{cl}w_{1}&\mbox{ if }\ |w_{2}|\leq|w_{1}|,\\ w_{2}&\mbox{ if }\ |w_{1}|<|w_{2}|,\end{array}\right.$
and
$g_{t+1}(w_{1},\ldots,w_{t+1}):=g_{2}(g_{t}(w,\ldots,t_{t}),w_{t+1}).$
For the proofs of both theorems under study, we shall proceed by induction with respect to the dimension $k=\dim A$. The case $k=0$ is straightforward by the above lemmas.
So assume that the conclusion holds for the subsets of $K^{n}$ of dimension $<k$ with $1\leq k\leq n$. We shall first consider the case $k<n$, reducing the problem to the sets $A$ of a special form. To this end, we shall apply parametrized cell decomposition, Lemma 2.1 and a model-theoretic compactness argument.
In order to prove Theorem 1.1, we can and shall additionally assume that the subset $A$ is bounded. Indeed, put
$A=\bigcup\,\{A_{I}:\ I\subset\{1,\ldots,n\}\}\ \ \text{with}\ \ A_{I}:=A\cap B_{I}$
with
$B_{I}:=\{x\in K^{n}:\ |x_{i}|\leq 1\ \ \text{for}\ i\in I\ \ \text{and}\ \ |x_{j}|>1\ \ \text{for}\ j\not\in I\},$
and
$\phi_{I}:B_{I}\to K^{n},\ \ \ \phi_{I}(x)=y,$
where $y_{i}=x_{i}$ for $i\in I$ and $y_{j}=1/x_{j}$ for $j\not\in I$. Then $B_{I}$ are clopen subsets of $K^{n}$, $\phi_{I}$ are homeomorphisms of $B_{I}$ onto the bounded subsets $D_{I}:=\phi_{I}(B_{I})$, and $E_{I}:=\phi_{I}(A_{I})$ are closed subsets of $D_{I}$. Further, the (topological) closures $\overline{E_{I}}$ of the sets $E_{I}$ are bounded subsets of $K^{n}$. Therefore, if $\overline{E_{I}}$ is the zero locus of a continuous 0-definable function $g_{I}:K^{n}\to K$, then $A_{I}$ is the zero locus of the function $f_{I}:=g_{I}\circ\phi_{I}$ which is a continuous 0-definable function on the clopen subset $B_{I}$. This establishes the reduction.
For coordinates $x=(x_{1},\ldots,x_{n})$ in the affine space $K^{n}$, write
$x=(y,z),\ \ y=(x_{1},\ldots,x_{k}),\ \ z=(x_{k+1},\ldots,x_{n}).$
Let $\pi:K^{n}\to K^{k}$ be the projection onto the first $k$ coordinates. For $y\in K^{k}$, denote by $A_{y}\subset K^{n-k}$ the fiber of the set $A$ over the point $y$.
Denote by $\overline{E}$ and $\partial E:=\overline{E}\setminus E$ the closure and frontier (in the valuation topology) of a set $E$, respectively. We have the basic dimension inequality (cf. [4, Proposition 5.3.4]):
(2.1) $\dim\partial E<\dim E.$
By Lemma 2.1, along with parametrized cell decomposition (op.cit. Theorem 5.7.3) and the induction hypothesis, we can assume that $A$ is (perhaps after a permutation of variables) the closure of a parametrized 0-definable cell $C=(C_{\xi})_{\xi}$ of dimension $k$ of cell-type $(1,\ldots,1,0,\ldots,0)$, with $RV$-sort parameters $\xi$ and centers $c_{\xi}$. Since definable $RV$-unions of finite sets stay finite (op.cit., Corollary 2.6.7), the restriction of $\pi$ to $C$ has finite fibers of bounded cardinality.
Further, by suitable 0-definable partitioning, we can assume that $A$ is the closure $\overline{E}$ of a 0-definable subset $E$ of dimension $k$ such that all the fibers $E_{y}$, $y\in\pi(E)$, have the same cardinality, say $s$, and the sets
$C_{j}(y)=\{c_{ji}(y),\ i=1,\ldots,s_{j}\}$
of the $j$-th coordinates of points from the fibers $E_{y}$ have the same cardinality, say $s_{j}$, for each $j=k+1,\ldots,n$. Since the fibers $E_{y}$ are finite, the projection
$F:=\pi(E)\subset K^{k}$
is of dimension $k$ (op.cit., Corollary 5.3.4). Again, this fact, along with cell decomposition, Lemma 2.1 and the induction hypothesis, allow us to come down to the case where $F$ is an open 0-definable subset of $K^{k}$.
Now consider the polynomials
$P_{j}(y,Z_{j}):=\prod_{z\in C_{j}(y)}\,(Z_{j}-z)=\prod_{i=1}^{s_{j}}\,(Z_{j}-c_{ji}(y)),\ \ y\in F,\ j=k+1,\ldots,n,$
Then
$P_{j}(y,Z_{j})=Z_{j}^{s_{j}}+b_{j,1}(y)Z_{j}^{s_{j-1}}+\ldots+b_{j,s_{j}}(y),\ \ j=k+1,\ldots,n,$
where $b_{j,i}:F\to K$, $i=1,\ldots,s_{j}$, are 0-definable functions.
We still need the following lemma, which resembles to some extent the primitive element theorem from algebraic geometry.
Lemma 2.3.
There exist a finite number of linear functions
$\lambda_{l}:K^{n-k}\to K,\ \ l=1,\ldots,p,$
with integer coefficients such that, for every $y\in F$, $\lambda_{l}$ is injective on the Cartesian product $\prod_{j=k+1}^{n}\,C_{j}(y)$ for some $l=1,\ldots,p$.
Proof.
The conclusion follows by a routine model-theoretic compactness argument. ∎
By Lemma 2.3 and Lemma 2.1, we can further assume, after a suitable 0-definable partitioning, that one linear function
$\lambda:K^{n-k}\to K$
with integer coefficients is injective on every product
$\prod_{j=k+1}^{n}\,C_{j}(y),\ \ y\in F.$
Consider now the polynomial
$P(y,Z):=\prod_{z\in E_{y}}\,(Z-\lambda(z))=Z^{s}+b_{1}(y)Z^{s-1}+\ldots+b_{s}(y),$
where $b_{j}:F\to K$ are 0-definable functions. Then
$E=\{x=(y,z)\in F\times K^{n-k}:$
$P_{k+1}(x_{1},\ldots,x_{k},x_{k+1})=\ldots=P_{n}(x_{1},\ldots,x_{k},x_{n})=$
$P(x_{1},\ldots,x_{k},\lambda(x_{k+1},\ldots,x_{n}))=0\}.$
The sets of all points at which the functions $b_{ji}(y)$ and $b_{i}(y)$ are not continuous are 0-definable subsets of $F$ of dimension $<k$ (op.cit., Theorem 5.1.1), and so are the closures of those sets by inequality 2.1. Hence and by Lemma 2.1 along with the induction hypothesis, we can additionally assume that $b_{ji}(y)$ and $b_{i}(y)$ are continuous functions on the open subset $F$. Then $E$ is a closed subset of $F\times K^{n-k}$, and thus
(2.2) $\partial E\subset\partial F\times K^{n-k}.$
In this manner, we have reduced the proofs of the theorems under study to the case where $A$ is the closure of the set $E$ described above. Moreover, in the proof of the first theorem, we have assumed without loss of generality that the set $E$ is bounded.
Proof of Theorem 1.1. Since $F$ is an open subset of $K^{k}$, its frontier $\partial F$ is a closed subset of $K^{k}$ of dimension $<k$ (inequality 2.1). By the induction hypothesis, $\partial F$ is the zero locus of a continuous 0-definable function $f:K^{k}\to K$.
Observe now that the functions $b_{ji}(y)$ are bounded because so are the sets $A$ and $E$ under study. Therefore the functions
$f(y)\cdot b_{ji}(y)\ \ \text{and}\ \ f(y)\cdot b_{i}(y)$
extend by zero through $\partial F$ to continuous functions on $\overline{F}$. And then they extend by zero off $F$ to continuous 0-definable functions on $K^{k}$.
We can thus regard the coefficients of the following polynomials (in the indeterminates $Z_{k+1},\ldots,Z_{n}$ and $Z$, respectively):
$Q_{k+1}(y,Z_{j}):=f(y)\cdot P_{k+1}(y,Z_{k+1}),\ \ldots\ ,\ Q_{n}(y,Z_{n}):=f(y)\cdot P_{n}(y,Z_{n})$
and
$Q(y,Z):=f(y)\cdot P(y,Z),$
as continuous 0-definable functions on $K^{k}$ vanishing off the subset $F$. Put
$G:=\{x\in K^{n}:\ Q_{k+1}(x_{1},\ldots,x_{k},x_{k+1})=\ldots=Q_{n}(x_{1},\ldots,x_{k},x_{n})=$
$Q(x_{1},\ldots,x_{k},\lambda(x_{k+1},\ldots,x_{n}))=0\}.$
Then
(2.3) $G\cap(F\times K^{n-k})=E$
and
(2.4) $G\cap((K^{k}\setminus F)\times K^{n-k})=(K^{k}\setminus F)\times K^{n-k}.$
Put
:=$\mathcal{E}:=\left\{(b,c,z)\in E\times K^{n-k}:\ (b,c)\in E\ \wedge\ \forall\,y\in\partial F\ |z|<|y-b|\right.$
$\left.\wedge\ \ \forall\,v,w\in K^{n-k}\ [((b,v),(b,w)\in E,\ v\neq w)\Rightarrow|z|<|v-w|]\right\}$
and
$\widetilde{E}:=p(\mathcal{E})\subset K^{n}=K^{k}\times K^{n-k},$
where
$p:K^{k}\times K^{n-k}\times K^{n-k}\ni(y,z,w)\mapsto(y,z+w)\in K^{k}\times K^{n-k}.$
Let $\widetilde{A}$ be the closure of $\widetilde{E}$; obviously, $E\subset\widetilde{E}$ and $A=\overline{E}\subset\widetilde{A}$.
Remark 2.4.
Note that the third condition of the conjunction in the definition of the set $\mathcal{E}$ will be used only in the proof of Theorem 1.2.
It is easy to check that $\widetilde{E}\supset E$ is a clopen subset of $F\times K^{n-k}$. In view of equality 2.2 and the second condition of the conjunction in the definition of the set $\mathcal{E}$, we get
$\partial\widetilde{E}=\partial E,\ \ \widetilde{A}=\widetilde{E}\cup\partial E$
and
$\widetilde{A}\cap((K^{k}\setminus F)\times K^{n-k})=\widetilde{A}\cap(\partial F\times K^{n-k})=\partial E.$
Hence and by eqaulity 2.4, we get
$G\cap\widetilde{A}\cap((K^{k}\setminus F)\times K^{n-k})=\widetilde{A}\cap((K^{k}\setminus F)\times K^{n-k})=\partial E.$
Further, equality 2.3 yields
$G\cap\widetilde{A}\cap(F\times K^{n-k})=E\cap\widetilde{A}=E.$
Therefore, combining the above two formulae, we get
(2.5) $G\cap\widetilde{A}=E\cup\partial E=\overline{E}=A.$
But by the induction hypothesis, $\partial\widetilde{E}=\partial E$ is the zero locus of a continuous 0-definable function $e:K^{n}\to K$. Then the function
$\widetilde{e}(x)=\left\{\begin{array}[]{cl}0&\mbox{ if }\ x\in\widetilde{A},\\ e(x)&\mbox{ if }\ x\in K^{n}\setminus\widetilde{A}.\end{array}\right.$
is continuous with zero locus $\widetilde{A}$:
$\widetilde{A}=\{x\in K^{n}:\ \widetilde{e}(x)=0\}.$
Hence and by equality 2.5, we obtain
(2.6) $A=\{Q_{k+1}(x_{1},\ldots,x_{k},x_{k+1})=\ldots=Q_{n}(x_{1},\ldots,x_{k},x_{n})=$
$Q(x_{1},\ldots,x_{k},\lambda(x_{k+1},\ldots,x_{n}))=\widetilde{e}(x)=0\}\subset K^{n}.$
Now it follows immediately from Lemma 2.2 that $A$ is the zero locus of the continuous 0-definable function
$g_{n-k+2}(Q_{k+1}(x),\ldots,Q_{n}(x),Q(x),\widetilde{e}(x)),$
as desired.
Finally, suppose that $A$ is of dimension $k=n$. Then $A=U\cup E$ for an open 0-definable subset $U\subset K^{n}$ and a closed 0-definable subset $E\subset K^{n}$ of dimension $<n$. By the induction hypothesis, $E$ is the zero locus of a continuous 0-definable function $f:K^{n}\to K$. Then $A$ is the zero locus of the following continuous definable function
$g(x)=\left\{\begin{array}[]{cl}0&\mbox{ if }\ x\in A,\\ f(x)&\mbox{ if }\ x\in K^{n}\setminus A.\end{array}\right.$
This completes the proof of Theorem 1.1. $\Box$
Proof of Theorem 1.2. We keep the notation from the foregoing proof and assume that the conclusion holds for the subsets of $K^{n}$ of dimension $<k$ with $1\leq k<n$. Again, we are reduced to the case where the set $A$ is the closure $\overline{E}$ of the set $E$ described before. The third condition of the conjunction in the definition of the set $\mathcal{E}$ ensures the following property of the set $\widetilde{E}$:
For each $x=(y,z)\in\widetilde{E}$ there is a unique $(y,w)\in E$, denoted by $\rho(x)$, such that
$v(z-w)=\min\,\{v(z-u):\ (y,u)\in E\}.$
But, by the induction hypothesis, there exists a 0-definable retraction $p:K^{n}\to\partial E$. Then the map $r:K^{n}\to A$ given by the formula
$r(x):=\left\{\begin{array}[]{cl}\rho(x)&\mbox{ if }\ x\in\widetilde{E},\\ p(x)&\mbox{ if }\ x\in K^{n}\setminus\widetilde{E}\end{array}\right.$
is a 0-definable retraction we are looking for.
Finally, suppose that $A$ is of dimension $k=n$ and the conclusion holds for the subsets of $K^{n}$ of dimension $<n$. As before, $A=U\cup E$ for an open 0-definable subset $U\subset K^{n}$ and a closed 0-definable subset $E\subset K^{n}$ of dimension $<n$. By the induction hypothesis, there exists a 0-definable retraction $p:K^{n}\to E$. Then the map $r:K^{n}\to A$ given by the formula
$r(x):=\left\{\begin{array}[]{cl}x&\mbox{ if }\ x\in U,\\ p(x)&\mbox{ if }\ x\in K^{n}\setminus U,\end{array}\right.$
is a 0-definable retraction we are looking for. This finishes the proof of Theorem 1.2. $\Box$
We conclude with the following comment. In the purely topological setting, the existence of a continuous retraction onto any non-empty closed subset of an ultranormal metrizable space was established by Dancis [7]. Previously, Ellis [9, 10] studied the problem of extension of continuous maps defined on closed subsets of zero-dimensional spaces with values in various types of metric spaces; in particular, of continuous functions from ultranormal and ultraparacompact spaces into a complete field with non-Archimedean absolute value. He obtained, among others, a non-Archimedean analogue of the Tietze–Urysohn theorem on extension of continuous functions from a closed subset of an ultranormal space into a locally compact field with non-Archimedean absolute value.